# Theory of Valley Splitting in Si/SiGe Spin-Qubits: Interplay of Strain, Resonances and Random Alloy Disorder

### v2

@@ A.1 Lattice-Periodic Part
$\sum_{n^{\prime}}\langle\Psi_{n}|\hat{H}_{\mathrm{cr}}|\Psi_{n^{\prime}}\rangle$ $=V\sum_{\nu}\left|\alpha_{\nu}\right|^{2}\sum_{\mathbf{k}}F_{n,\nu}^{*}\left(\mathbf{k}\right)\times$
@@
where we used the strong localization of the valley-envelopes in reciprocal space such that
$F_{n,\nu^{\prime}}\left(\mathbf{k}+\mathbf{k}_{\nu}-\mathbf{k}_{\nu^{\prime}}\right)\approx\delta_{\nu,\nu^{\prime}}F_{n,\nu}\left(\mathbf{k}\right).$
Finally, inverse Fourier transform
$F_{n,\nu}\left(\mathbf{k}\right)=\frac{1}{V}\int_{V}\mathrm{d}^{3}r\,\mathrm{e}^{-i\mathbf{k}\cdot\mathbf{r}}F_{n,\nu}\left(\mathbf{r}\right)$ (52)
and integration by parts yields
$\sum_{n^{\prime}}$ $\langle\Psi_{n}|\hat{H}_{\mathrm{cr}}|\Psi_{n^{\prime}}\rangle=$ (53)
$=\sum_{\nu}\left|\alpha_{\nu}\right|^{2}\frac{\hbar^{2}}{2}\mathbf{n}\cdot m_{n,\mathbf{k}_{\nu}}^{-1}\bigg{(}F_{n,\nu}^{*}\left(\mathbf{r}\right)\nabla F_{n,\nu}\left(\mathbf{r}\right)-$
[…]
$\phantom{=}+\sum_{\nu}\left|\alpha_{\nu}\right|^{2}\int_{V}\mathrm{d}^{3}r\,F_{n,\nu}^{*}\left(\mathbf{r}\right)\bigg{(}\left(E_{n,\mathbf{k}_{\nu}}-E_{n}\right)F_{n,\nu}\left(\mathbf{r}\right)$
[…]
The first term is a boundary condition for the envelope wave functions. The second term includes the band edge energy and the kinetic energy in effective mass approximation for the coupled envelope equation model.
A.2 Heterostructure Part
In the following, the corresponding computation is carried out for the heterostructure part, which finally yields a coupling between the different valley-specific envelope wave functions. The ansatz (48) with (50) yields
$\sum_{n^{\prime}}$ $\langle\Psi_{n}|\hat{U}|\Psi_{n^{\prime}}\rangle=\sum_{n^{\prime}}\sum_{\mathbf{k},\mathbf{k}^{\prime}}F_{n,\mathbf{k}}^{*}F_{n^{\prime},\mathbf{k}^{\prime}}\times$
×$\phantom{=}\times\int_{V}\mathrm{d}^{3}r\,\mathrm{e}^{-i\left(\mathbf{k}-\mathbf{k}^{\prime}\right)\cdot\mathbf{r}}\,U\left(\mathbf{r}\right)u_{n,\mathbf{k}}^{*}\left(\mathbf{r}\right)u_{n^{\prime},\mathbf{k}^{\prime}}\left(\mathbf{r}\right)$
$=$ $\sum_{n^{\prime}}\sum_{\mathbf{k},\mathbf{k}^{\prime}}\sum_{\nu,\nu^{\prime}}\alpha_{\nu}^{*}\alpha_{\nu^{\prime}}F_{n,\nu}^{*}\left(\mathbf{k}\right)F_{n^{\prime},\nu^{\prime}}\left(\mathbf{k}^{\prime}\right)\int_{V}\mathrm{d}^{3}r\,\times$
×$\times\mathrm{e}^{-i\left(\mathbf{k}-\mathbf{k}^{\prime}+\mathbf{k}_{\nu}-\mathbf{k}_{\nu^{\prime}}\right)\cdot\mathbf{r}}\,U\left(\mathbf{r}\right)u_{n,\mathbf{k}_{\nu}+\mathbf{k}}^{*}\left(\mathbf{r}\right)u_{n^{\prime},\mathbf{k}_{\nu^{\prime}}+\mathbf{k}^{\prime}}\left(\mathbf{r}\right).$
The dispersion of the near band-edge Bloch factors is ignored by using the approximation
$u_{n,\mathbf{k}_{\nu}+\mathbf{k}}\left(\mathbf{r}\right)$ $\approx u_{n,\mathbf{k}_{\nu}}\left(\mathbf{r}\right).$
With this and the inverse Fourier transform (52) we arrive at
$\sum_{n^{\prime}}\langle\Psi_{n}|\hat{U}$ $|\Psi_{n^{\prime}}\rangle\approx\sum_{n^{\prime}}\sum_{\nu,\nu^{\prime}}\alpha_{\nu}^{*}\alpha_{\nu^{\prime}}\int_{V}\mathrm{d}^{3}r\,\mathrm{e}^{-i\left(\mathbf{k}_{\nu}-\mathbf{k}_{\nu^{\prime}}\right)\cdot\mathbf{r}}\times$
$\times$ $F_{n,\nu}^{*}\left(\mathbf{r}\right)F_{n^{\prime},\nu^{\prime}}\left(\mathbf{r}\right)U\left(\mathbf{r}\right)u_{n,\mathbf{k}_{\nu}}^{*}\left(\mathbf{r}\right)u_{n^{\prime},\mathbf{k}_{\nu^{\prime}}}\left(\mathbf{r}\right).$
This expression is separated into the intravalley ($\nu=\nu^{\prime}$) and intervalley ($\nu\neq\nu^{\prime}$) contributions. Since the Bloch factors vary on an atomistic scale, i.e., much faster than the confinement potential and the envelope wave function, the intravalley part is obtained as [67]
$\sum_{n^{\prime}}$ $\langle\Psi_{n}|\hat{U}|\Psi_{n^{\prime}}\rangle\bigg{|}_{\mathrm{intra}}=\sum_{n^{\prime}}\sum_{\nu}\alpha_{\nu}^{*}\alpha_{\nu}\int_{V}\mathrm{d}^{3}r\,\times$
×$\phantom{\approx}\times F_{n,\nu}^{*}\left(\mathbf{r}\right)F_{n^{\prime},\nu}\left(\mathbf{r}\right)U\left(\mathbf{r}\right)u_{n,\mathbf{k}_{\nu}}^{*}\left(\mathbf{r}\right)u_{n^{\prime},\mathbf{k}_{\nu}}\left(\mathbf{r}\right)$
$\approx\sum_{n^{\prime}}\sum_{\nu}\left|\alpha_{\nu}\right|^{2}\Omega_{p}\sum_{j=1}^{N}F_{n,\nu}^{*}\left(\mathbf{R}_{j}\right)F_{n^{\prime},\nu}\left(\mathbf{R}_{j}\right)U\left(\mathbf{R}_{j}\right)$
@@ A.2 Heterostructure Part
×$\phantom{\approx}\times\frac{1}{\Omega_{p}}\int_{\Omega_{p}}\mathrm{d}^{3}r\,u_{n,\mathbf{k}_{\nu}}^{*}\left(\mathbf{r}\right)u_{n^{\prime},\mathbf{k}_{\nu}}\left(\mathbf{r}\right)$
$\approx\sum_{\nu}\left|\alpha_{\nu}\right|^{2}\int_{V}\mathrm{d}^{3}r\,F_{n,\nu}^{*}\left(\mathbf{r}\right)U\left(\mathbf{r}\right)F_{n,\nu}\left(\mathbf{r}\right),$ (54)
where we used the orthogonality relation (47). The corresponding intervalley contribution is obtained as
$\sum_{n^{\prime}}$ $\langle\Psi_{n}|\hat{U}|\Psi_{n^{\prime}}\rangle\bigg{|}_{\mathrm{inter}}\approx$ (55)
$\approx$ $\sum_{n^{\prime}}\sum_{\nu,\nu^{\prime}\neq\nu}\alpha_{\nu}^{*}\alpha_{\nu^{\prime}}\sum_{\mathbf{G},\mathbf{G}^{\prime}}c_{n,\mathbf{k}_{\nu}}^{*}\left(\mathbf{G}\right)c_{n^{\prime},\mathbf{k}_{\nu^{\prime}}}\left(\mathbf{G}^{\prime}\right)\times$
×$\times\int_{V}\mathrm{d}^{3}r\,F_{n,\nu}^{*}\left(\mathbf{r}\right)\mathrm{e}^{-i\left(\mathbf{G}-\mathbf{G}^{\prime}+\mathbf{k}_{\nu}-\mathbf{k}_{\nu^{\prime}}\right)\cdot\mathbf{r}}U\left(\mathbf{r}\right)F_{n^{\prime},\nu^{\prime}}\left(\mathbf{r}\right).$
where we used a plane wave expansion of the Bloch factors.
A.3 Coupled Envelopes
The combination of Eqs. (53)–(55) yields a system of multi-valley coupled envelope equations
$E_{n}F_{n,\nu}\left(\mathbf{r}\right)$ $=-\frac{\hbar^{2}}{2}\nabla\cdot\left(m_{n,\mathbf{k}_{\nu}}^{-1}\nabla F_{n,\nu}\left(\mathbf{r}\right)\right)$
$\phantom{=}+\left(E_{n,\mathbf{k}_{\nu}}+U\left(\mathbf{r}\right)\right)F_{n,\nu}\left(\mathbf{r}\right)$
$\phantom{=}+\sum_{n^{\prime}}\sum_{\nu^{\prime}\neq\nu}\frac{\alpha_{\nu}^{*}\alpha_{\nu^{\prime}}}{\left|\alpha_{\nu}\right|^{2}}\sum_{\mathbf{G},\mathbf{G}^{\prime}}c_{n,\mathbf{k}_{\nu}}^{*}\left(\mathbf{G}\right)c_{n^{\prime},\mathbf{k}_{\nu^{\prime}}}\left(\mathbf{G}^{\prime}\right)\times$
[…]
In the case of a single band (conduction band only, band indices $n$, $n^{\prime}$ are suppressed in the following), two valleys at $\mathbf{k}_{\nu}=\pm\mathbf{k}_{0}$, symmetry factors $\alpha_{\nu}=1/\sqrt{2}$ and identical effective masses $m_{n,\mathbf{k}_{\nu}}\equiv m$ and band edge energies $E_{n,\mathbf{k}_{\nu}}\equiv E_{c}$, we obtain
@@ A.3 Coupled Envelopes
$\left(\begin{array}[]{cc}H_{0}\left(\mathbf{r}\right)&V_{c}\left(\mathbf{r}\right)\\ V_{c}^{*}\left(\mathbf{r}\right)&H_{0}\left(\mathbf{r}\right)\end{array}\right)\left(\begin{array}[]{c}F_{+}\left(\mathbf{r}\right)\\ F_{-}\left(\mathbf{r}\right)\end{array}\right)=E\left(\begin{array}[]{c}F_{+}\left(\mathbf{r}\right)\\ F_{-}\left(\mathbf{r}\right)\end{array}\right)$
with
$H_{0}\left(\mathbf{r}\right)$ $=-\frac{\hbar^{2}}{2}\nabla\cdot\left(m^{-1}\nabla\right)+E_{c}+U\left(\mathbf{r}\right),$
$V_{c}\left(\mathbf{r}\right)$ $=\sum_{\mathbf{G},\mathbf{G}^{\prime}}c_{+}^{*}\left(\mathbf{G}\right)c_{-}\left(\mathbf{G}^{\prime}\right)\mathrm{e}^{-i\left(\mathbf{G}-\mathbf{G}^{\prime}+2\mathbf{k}_{0}\right)\cdot\mathbf{r}}U\left(\mathbf{r}\right)$
Renaming $F_{\pm}\left(\mathbf{r}\right)\to\Psi_{\pm}\left(\mathbf{r}\right)$ and ignoring the constant band edge energy $E_{c}$ finally yields Eq. (1).
@@
Appendix B Evaluation of Valley Splitting Components and Band Structure Coefficients

### v3

@@ A.2 Multi-Valley Envelope Wave Function
$\Delta_{\mathbf{k}_{0}}\left(\mathbf{r}-\mathbf{r}^{\prime}\right)=\frac{1}{V}\sum_{\left(\mathbf{k}+\mathbf{k}_{0}\right)\in S\left(\mathbf{k}_{0}\right)}\mathrm{e}^{i\left(\mathbf{k}_{0}+\mathbf{k}\right)\cdot\left(\mathbf{r}-\mathbf{r}^{\prime}\right)}$ (44)
that leads to smoothing of the effective potential and ensures proper restriction to small wave vectors within the respective Brillouin zone sectors.
Using the identity for the lattice-periodic problem
$T_{n,n^{\prime}}\left(\mathbf{k}_{0}\right)+V_{n,n^{\prime}}\left(\mathbf{k}_{0}\right)+$ $\frac{\hbar}{m_{0}}\mathbf{k}_{0}\cdot\mathbf{p}_{n,n^{\prime}}\left(\mathbf{k}_{0}\right)=$
$=\left(E_{n,\mathbf{k}_{0}}-\frac{\hbar^{2}k_{0}^{2}}{2m_{0}}\right)\delta_{n,n^{\prime}},$
the system of coupled envelope equations takes the form
$Ef_{n,\mathbf{k}_{0}}\left(\mathbf{r}\right)$ $=\left(-\frac{\hbar^{2}}{2m_{0}}\nabla^{2}+E_{n,\mathbf{k}_{0}}\right)f_{n,\mathbf{k}_{0}}\left(\mathbf{r}\right)$ (45)
$-\frac{i\hbar}{m_{0}}\sum_{n^{\prime}}\left(\hbar\mathbf{k}_{0}\delta_{n,n^{\prime}}+\mathbf{p}_{n,n^{\prime}}\left(\mathbf{k}_{0}\right)\right)\cdot\nabla f_{n^{\prime},\mathbf{k}_{0}}\left(\mathbf{r}\right)$
$+\sum_{n^{\prime}}\sum_{\mathbf{k}_{0}^{\prime}}\int_{V}\mathrm{d}^{3}r^{\prime}\,U_{\mathbf{k}_{0},\mathbf{k}_{0}^{\prime}}^{n,n^{\prime}}\left(\mathbf{r},\mathbf{r}^{\prime}\right)f_{n^{\prime},\mathbf{k}_{0}^{\prime}}\left(\mathbf{r}^{\prime}\right).$
A.3 Elimination of Remote Bands
The multi-band system (45) can be reduced to an effective equation for the conduction band $n=c$ by eliminating remote bands $n\neq c$ (Löwdin renormalization). We approximate the remote band envelopes as [74]
$f_{n,\mathbf{k}_{0}}\left(\mathbf{r}\right)$ $\approx-\frac{i\hbar}{m_{0}}\frac{\mathbf{p}_{n,c}\left(\mathbf{k}_{0}\right)}{E_{c,\mathbf{k}_{0}}-E_{n,\mathbf{k}_{0}}}\cdot\nabla f_{c,\mathbf{k}_{0}}\left(\mathbf{r}\right)$
and neglect interaction between remote bands. Moreover, we ignore the renormalization of the confinement potential term for the conduction band. With this, we arrive at the single-band envelope equation in effective mass approximation
$Ef_{c,\mathbf{k}_{0}}\left(\mathbf{r}\right)$ $=-\frac{\hbar^{2}}{2}\nabla\cdot\left(m_{c,\mathbf{k}_{0}}^{-1}\nabla f_{c,\mathbf{k}_{0}}\left(\mathbf{r}\right)\right)+E_{c,\mathbf{k}_{0}}f_{c,\mathbf{k}_{0}}\left(\mathbf{r}\right)$
$\phantom{=}+\sum_{\mathbf{k}_{0}^{\prime}}\int_{V}\mathrm{d}^{3}r^{\prime}\,U_{\mathbf{k}_{0},\mathbf{k}_{0}^{\prime}}^{c,c}\left(\mathbf{r},\mathbf{r}^{\prime}\right)f_{c,\mathbf{k}_{0}^{\prime}}\left(\mathbf{r}^{\prime}\right)$ (46)
with the effective mass tensor
$m_{c,\mathbf{k}_{0}}^{-1}=\frac{1}{m_{0}}I+\frac{2}{m_{0}^{2}}\sum_{n\neq c}\frac{\mathbf{p}_{c,n}\left(\mathbf{k}_{0}\right)\otimes\mathbf{p}_{c,n}^{\dagger}\left(\mathbf{k}_{0}\right)}{E_{c,\mathbf{k}_{0}}-E_{n,\mathbf{k}_{0}}}.$ (47)
Here we have used the condition $\mathbf{p}_{c,c}\left(\mathbf{k}_{0}\right)=-\hbar\mathbf{k}_{0}$, which ensures a minimum of the conduction band at the position of the valley wave vector $\mathbf{k}_{0}$.
A.4 Local Approximation
If the momentum space envelopes are sufficiently localized near the valley states [75, 77], the low-pass filtered delta function (44) can be replaced by a Dirac delta leading to a local approximation of the envelope equation
$E$ $f_{c,\mathbf{k}_{0}}\left(\mathbf{r}\right)\approx$
$\approx$ $-\frac{\hbar^{2}}{2}\nabla\cdot\left(m_{c,\mathbf{k}_{0}}^{-1}\nabla f_{c,\mathbf{k}_{0}}\left(\mathbf{r}\right)\right)+\left(E_{c,\mathbf{k}_{0}}+U\left(\mathbf{r}\right)\right)f_{c,\mathbf{k}_{0}}\left(\mathbf{r}\right)$
$+\sum_{\mathbf{k}_{0}^{\prime}\neq\mathbf{k}_{0}}\mathrm{e}^{-i\left(\mathbf{k}_{0}-\mathbf{k}_{0}^{\prime}\right)\mathbf{r}}u_{\mathbf{k}_{0}}^{*}\left(\mathbf{r}\right)u_{\mathbf{k}_{0}^{\prime}}\left(\mathbf{r}\right)U\left(\mathbf{r}\right)f_{c,\mathbf{k}_{0}^{\prime}}\left(\mathbf{r}\right).$
Restriction to the two low-energy valleys $\mathbf{k}_{0}=\pm k_{0}\mathbf{e}_{z}$ in a biaxially strained [001] QW and renaming $f_{c,\pm\mathbf{k}_{0}}\left(\mathbf{r}\right)\to\Psi_{\pm}\left(\mathbf{r}\right)$ yields the coupled envelope model (1).
Appendix B Empirical Pseudopotential Model
The lattice-periodic part of the problem is modeled using an empirical pseudopotential model
$\frac{\hbar^{2}}{2m_{0}}\left\|\mathbf{G}+\mathbf{k}\right\|^{2}c_{n,\mathbf{k}}\left(\mathbf{G}\right)+$
$\hphantom{=}+\sum_{\mathbf{G}^{\prime}}V_{\mathrm{ps}}\left(\mathbf{G}+\mathbf{k},\mathbf{G}^{\prime}+\mathbf{k}\right)c_{n,\mathbf{k}}\left(\mathbf{G}^{\prime}\right)=E_{n,\mathbf{k}}c_{n,\mathbf{k}}\left(\mathbf{G}\right).$
The pseudopotential has a local and a nonlocal component
$V_{\mathrm{ps}}\left(\mathbf{K},\mathbf{K}^{\prime}\right)=V_{\mathrm{loc}}\left(\mathbf{K}-\mathbf{K}^{\prime}\right)+V_{\mathrm{nloc}}\left(\mathbf{K},\mathbf{K}^{\prime}\right),$
where $\mathbf{K}=\mathbf{G}+\mathbf{k}$ and $\mathbf{K}^{\prime}=\mathbf{G}+\mathbf{k}^{\prime}$. The local pseudopotential reads
$V_{\mathrm{loc}}\left(\mathbf{G}\right)$ $=S\left(\mathbf{G}\right)V_{\mathrm{atom}}\left(\left\|\mathbf{G}\right\|\right),$
where $V_{\mathrm{atom}}\left(\left\|\mathbf{G}\right\|\right)$ is the radially symmetric atomic pseudopotential for silicon and $\mathbf{G}$ is a reciprocal lattice vector of the fcc lattice. The diamond crystal structure is described via the structure factor
$S\left(\mathbf{G}\right)$ $=\cos\left(\frac{\mathbf{G}\cdot\boldsymbol{\mathbf{\tau}}}{2}\right),$
where $\boldsymbol{\tau}=a_{0}/4\times\left(1,1,1\right)^{T}$ is the offset between the two fcc sub-lattices. The nonlocal pseudopotential reads
$V_{\mathrm{nloc}}\left(\mathbf{K},\mathbf{K}^{\prime}\right)=\frac{4\pi R_{0}^{3}}{\Omega_{a}}A_{0}S\left(\mathbf{K}-\mathbf{K}^{\prime}\right)f_{0}\left(R_{0}K,R_{0}K^{\prime}\right),$
which describes a spherically symmetric $s$-well with depth $A_{0}$ and radius $R_{0}$. For a square well we use [52, 81]
$f_{0}\left(x,x^{\prime}\right)$ $=\frac{xj_{1}\left(x\right)j_{0}\left(x^{\prime}\right)-x^{\prime}j_{1}\left(x^{\prime}\right)j_{0}\left(x\right)}{\left(x+x^{\prime}\right)\left(x-x^{\prime}\right)},$
where $j_{\nu}\left(x\right)$ are the spherical Bessel functions. Following Ref. [54], spin-orbit interaction is ignored for silicon.
symbol description value
$E_{\mathrm{cutoff}}$ cutoff energy $12.0\,\mathrm{Ry}$
$V_{\mathrm{atom}}\left(\sqrt{3}\right)$ atomic pseudopotentia shape parameters … $-0.2241\;\mathrm{Ry}$ [54]
$V_{\mathrm{atom}}\left(\sqrt{8}\right)$ $\phantom{+}0.0520\;\mathrm{Ry}$ [54]
$V_{\mathrm{atom}}\left(\sqrt{11}\right)$ $\phantom{+}0.0724\;\mathrm{Ry}$ [54]
$A_{0}$ s-well depth $0.03\;\mathrm{Ry}$ [54]
@@ Appendix B Empirical Pseudopotential Model
$k_{F}$ Fermi wave number $1.66\times 2\pi/a_{0}$ [55]
$\zeta$ internal strain parameter 0.530.530.53 [54]
$a_{0}$ lattice constant $0.543\,\mathrm{nm}$
Tab. 3: Parameters of the nonlocal empirical pseudopotential model for silicon.
B.0.1 Strain
Strain leads to a deformation of the crystal and thereby to a modification of the primitive lattice vectors
$\mathbf{a}_{i}^{\prime}$ $=\left(I+\varepsilon\right)\mathbf{a}_{i},$
where the relaxed basis vectors of the fcc lattice are
$\mathbf{a}_{1}$ $=\frac{a_{0}}{2}\left(\begin{array}[]{c}0\\ 1\\ 1\end{array}\right),$ $\mathbf{a}_{2}$ $=\frac{a_{0}}{2}\left(\begin{array}[]{c}1\\ 0\\ 1\end{array}\right),$ $\mathbf{a}_{3}$ $=\frac{a_{0}}{2}\left(\begin{array}[]{c}1\\ 1\\ 0\end{array}\right).$
As a consequence, the basis vectors of the reciprocal lattice are modified according to
@@ B.0.1 Strain
$\mathbf{b}_{i}^{\prime}=\frac{\pi}{\Omega_{p}^{\prime}}\sum_{j,k}\epsilon_{i,j,k}\left(\mathbf{a}_{j}^{\prime}\times\mathbf{a}_{k}^{\prime}\right)$
where $\Omega_{p}^{\prime}=\mathbf{a}_{1}^{\prime}\cdot\left(\mathbf{a}_{2}^{\prime}\times\mathbf{a}_{3}^{\prime}\right)$ is the volume of the strained primitive unit cell and $\epsilon_{i,j,k}$ is the totally antisymmetric Levi–Civita tensor. In the small strain limit, the modified reciprocal lattice vectors are well approximated by $\mathbf{b}_{i}^{\prime}\approx\left(I-\varepsilon\right)\mathbf{b}_{i}$. The reciprocal lattice vectors of the strained crystal are integer multiples of the basis vectors $\mathbf{G}^{\prime}=\sum_{i=1}^{3}n_{i}\mathbf{b}_{i}^{\prime}$. The set of reciprocal lattice vectors is truncated at $\left\|\mathbf{G}\right\|^{2}\leq 2m_{0}E_{\mathrm{cutoff}}/\hbar^{2}$. In the main text, the prime to indicate strain is omitted.
B.0.2 Interpolation of the Atomic Pseudopotential
Because of the strain-induced modification of the reciprocal lattice vectors, the atomic pseudopotential $V_{\mathrm{atom}}\left(G\right)$ is required at all possible values of $G$. We follow the usual approach [54, 55, 57] and employ a cubic spline interpolation using the form factors at $G=\left\{\text{$\sqrt{3}$},\text{$\sqrt{8}$},\text{$\sqrt{11}$}\right\}\,2\pi/a_{0}$ and boundary conditions $V_{\mathrm{atom}}\left(0\right)=-\frac{2}{3}E_{F}$, $V_{\mathrm{atom}}^{\prime}\left(0\right)=0$, $V_{\mathrm{atom}}\left(3k_{F}\right)=0$ and $V_{\mathrm{atom}}^{\prime}\left(3k_{F}\right)=0$. The Fermi energy is taken as $E_{F}=\hbar^{2}k_{F}^{2}/\left(2m_{0}\right).$ Parameter values are listed in Tab. B.
B.0.3 Internal Ionic Displacement
In addition to the deformation of the unit cell, we consider an additional displacement of the atoms arising from bond-length distortion and bond-bending [55, 57]. Bond-length distortion results from the (macroscopic) deformation of the regular tetrahedral structure, where ions tend to shift towards the barycenter of the distorted tetrahedron
$\boldsymbol{\tau}_{0}$ $=\left(I+\varepsilon\right)\boldsymbol{\tau}.$
This ionic displacement is accompanied by bond-bending and electrostatic repulsion from nearest neighbors which tend to restore its original position at the circumcenter of the tetrahedron. For small strain this is
@@ B.0.3 Internal Ionic Displacement
$\boldsymbol{\tau}_{1}$ $\approx\left(I+\varepsilon\right)\boldsymbol{\tau}-\frac{a_{0}}{2}\left(\begin{array}[]{c}\varepsilon_{y,z}\\ \varepsilon_{z,x}\\ \varepsilon_{x,y}\end{array}\right),$
which deviates from the barycenter in the presence of shear strain. As a consequence, the ions take a position between the circumcenter $\boldsymbol{\tau}_{0}$ and the barycenter $\boldsymbol{\tau}_{1}$ of the distorted tetrahedron. The ionic displacement is modeled phenomenologically as
$\boldsymbol{\tau}^{\prime}$ $=\left(1-\zeta\right)\boldsymbol{\tau}_{0}+\zeta\boldsymbol{\tau}_{1}\approx\left(I+\varepsilon\right)\boldsymbol{\tau}-\frac{a_{0}}{2}\left(\begin{array}[]{c}\varepsilon_{y,z}\\ \varepsilon_{z,x}\\ \varepsilon_{x,y}\end{array}\right)\zeta,$
where $\zeta$ is the internal strain parameter, cf. Fig. 1 (e).
@@
Appendix C Evaluation of Valley Splitting Components and Bloch Factor Coefficients